# Stochastic Deep Learning Surrogate Models for Uncertainty Propagation in Microstructure-Properties of Ceramic Aerogels

### v1

@@ Abstract
Deep learning surrogate models have become pivotal in enabling model-driven materials discovery to achieve exceptional properties. However, ensuring the accuracy and reliability of predictions from these models, trained on limited and sparse material datasets remains a significant challenge. This study introduces an integrated deep learning framework for predicting the synthesis, microstructure, and mechanical properties of ceramic aerogels, leveraging physics-based models such as Lattice Boltzmann simulations for microstructure formation and stochastic finite element methods for mechanical property calculations. To address the computational demands of repeated physics-based simulations required for experimental calibration and material design, a linked surrogate model is developed, leveraging Convolutional Neural Networks (CNNs) for stochastic microstructure generation and microstructure-to-mechanical property mapping. To overcome challenges associated with limited training datasets from expensive physical modeling, CNN training is formulated within a Bayesian inference framework, enabling robust uncertainty quantification in predictions. Numerical results highlight the strengths and limitations of the linked surrogate framework, demonstrating its effectiveness in predicting properties of aerogels with pore sizes and morphologies similar to the training data (in-distribution) and its ability to interpolate to new microstructural features between training data (out-of-distribution).
@@
Bayesian inference , convolutional neural networks , surrogate modeling , uncertainty quantification , ceramic aerogel
@@ 1 Introduction
Ceramic aerogels, composed of interconnected amorphous ceramic nanoparticles, exhibit a highly porous structure and exceptional thermal insulation properties, making them ideal for advanced thermal management in extreme environments. Their applications span civil and military aviation [1], space exploration [2], fire-resistant wearables [3], microelectronics [4, 5], and energy-efficient building systems [6, 7]. However, widespread adoption is hindered by their low mechanical stability. Overcoming this requires model-driven materials discovery [8, 9, 10] to predict synthesis-microstructure-property relationships for optimizing synthesis parameters to create mechanically robust aerogels without compromising insulation performance.
@@
Addressing computational challenges in uncertainty quantification, inference, and optimization for model-driven materials discovery hinges on developing surrogate models that approximate synthesis-to-property relationships. Trained on high-fidelity data, these models enable faster and more efficient predictions. Convolutional Neural Networks (CNNs) have become prominent in computational materials science [11, 12, 13, 14, 15], in generating realistic microstructural images and predicting macroscopic properties. However, their deployment in high-consequence applications, such as materials discovery, faces challenges, including sparse, limited, and stochastic training datasets from physical simulations. Traditional training methods relying on maximum likelihood estimation are prone to overfitting, producing overconfident and unreliable predictions. Furthermore, current approaches for ensuring CNN reliability depend on large datasets, limiting their effectiveness in small-data regimes. Framing training as a Bayesian inference problem offers a robust solution to handling small and uncertain datasets in deep learning surrogate modeling for physical simulations [16, 17]. This approach represents parameter uncertainty through probability distributions, reducing overfitting and enabling uncertainty quantification [18, 19]. However, applying Bayesian inference to CNNs is computationally challenging due to the large number of weight parameters. Zhu et al. [20] addressed this by employing a variational gradient descent algorithm based on Stein’s method, extending Bayesian inference to convolutional encoder-decoder networks. Their approach captured uncertainty in flow through heterogeneous media, even in scenarios where the input (permeability) and output (flow and pressure) fields lacked shared underlying structures. Similarly, Shridhar et al. [21] applied variational inference to estimate posterior weight distributions in CNNs, effectively propagating uncertainty in classification tasks across datasets like MNIST and CIFAR-100.
This work introduces an integrated modeling framework (Figure 1) for simulating the synthesis, microstructure, and properties of ceramic aerogels with high computational efficiency. Specifically, it focuses on a novel synthesis method utilizing in situ bubble-supported pore formation with foaming agents [22, 3], which addresses the limitations of conventional supercritical drying by preserving the porous structure without collapsing the solid matrix. This synthesis enables ambient-pressure, room-temperature drying, significantly reducing processing time while facilitating scalable and cost-effective production of ceramic aerogels. The synthesis-microstructure physics-based model in Figure 1, relies on Lattice Boltzmann simulations of the foaming process, capturing stochastic bubble nucleation, bubble dynamics, and coalescence to predict microstructure formation. The microstructure-property model employs stochastic finite element simulations to analyze the elastic deformation of the solid phase of the aerogel. Using the microstructural patterns, this model computes the displacement and strain fields to estimate the mechanical properties of the aerogel, i.e., the strain energy as the quantity of interest (QoI) for the modeling framework. To enable efficient simulations of synthesis, microstructure, and properties, we develop linked surrogates that demonstrate superior representations of the coupled system compared to composite surrogate models directly mapping synthesis inputs to QoI [23]. The microstructure surrogate leverages a scalable generative network trained on microscopic images, enabling the generation of morphologically consistent microstructures across larger domains necessary for macroscopic property predictions. Additionally, the surrogate model of the material property employs a CNN to map microstructural images to the QoI. A key innovation of this framework is the formulation of CNN training as Bayesian inference, enabling the estimation of the probability distribution of weight parameters using high-fidelity training data. To address the computational challenges of Bayesian inference in high-dimensional parameter spaces, we extend the Variational Inference approach used in Bayesian Neural Networks to accommodate the complexity of CNN architectures. The resulting Bayesian Convolutional Neural Network (BayesCNN) mitigates overconfidence in parameter estimation, supports training on small and uncertain datasets derived from high-fidelity simulations, and quantifies the uncertainty in surrogate model predictions. The linked surrogates are systematically employed to characterize and propagate various sources of uncertainty through the synthesis-microstructure-property relationship and the computational prediction of the QoI. This includes capturing microstructural randomness and quantifying errors introduced during surrogate model construction.
@@
Figure 1: Flowchart of the integrated modeling framework developed in this work for propagating uncertainty through synthesis inputs $\boldsymbol{\xi}$, microstructure spatial field $\varphi(\mathbf{x})$, and mechanical property $Q$ of ceramic aerogels. The physics-based models, $\mathcal{F}_{LB}(\cdot)$ and $\mathcal{F}_{FE}(\cdot)$, are parameterized by $\boldsymbol{\theta}_{LB}$ and $\boldsymbol{\theta}_{FE}$, and utilize synthesis inputs and microstructural patterns to predict material properties. Computational efficiency is achieved through surrogate models $\mathcal{G}_{GAN}(\cdot)$ and $\mathcal{G}_{CNN}(\cdot)$, constructed using high-fidelity training data from physical simulations. The parameter of the surrogate models include the stochastic latent vector $\boldsymbol{z}$, deterministic weights of the generative network $\boldsymbol{w}_{GAN}$, and posterior probability distributions of the wights $\boldsymbol{w}_{CNN}$ from the Bayesian inference.
@@
The remainder of this manuscript is structured as follows. Section 2 provides an overview of the ceramic aerogel synthesis process, including its modeling using Lattice Boltzmann simulations and the development of surrogate models to represent microstructural patterns. Section 3 describes the stochastic finite element model used to establish microstructure-property relationships and presents the comprehensive formulation and efficient solution algorithm of BayesCNN for surrogate modeling. Section 4 outlines the construction of individual surrogate models, detailing the creation of training datasets from high-fidelity simulations and the validation of these models. Additionally, this section demonstrates the application of the linked-surrogates for propagating of uncertainty from synthesis parameters and stochastic microstructure to the QoI. Finally, Section 5 presents a discussion of the findings and the concluding remarks.
@@ 2.1 Lattice Boltzmann simulation of microstructure formation
where $\Delta t$ is the time step, $\tau$ is a relaxation time, and $f_{i}^{eq}$ is the distribution function at equilibrium [33]. The solution of Eq. (1) involves two iterative steps: streaming, where the distribution functions $f_{i}$ move to neighboring lattice sites, and collision, where $f_{i}$ is locally updated at each site. Macroscopic fluid properties, such as density and velocity, are then derived from the moments of $f_{i}$. Building on this foundation, FSLBM simulations utilize a fast mass-tracking algorithm [27] to monitor the interfaces between gas and liquid phases throughout the computational domain. The liquid-gas interface is modeled as a free surface, with the VOF method employed to track the interface using a level function, which represents the liquid fraction within each computational cell. Surface tension $\sigma_{LB}$ are incorporated by calculating the interface curvature from the level function and applying the resulting pressure jump across the interface. In this work, to capture the highly viscous and shear-thinning behavior of the aerogel precursor, governed by the concentration of viscosifier $C_{vis}$, the Carreau-Yasuda model [34] is adopted to describe the fluid’s rheological properties. The model provides an accurate representation of the non-Newtonian characteristics essential for simulating the foaming process. The influence of TEOS concentration on bubble growth is incorporated by modeling gas diffusion into the bubbles and within the liquid phase using the advection-diffusion equation for gas concentration $C$,
@@
Figure 2 presents snapshots of pre-aerogel foaming simulations conducted using FSLBM, illustrating the effects of varying foaming agent concentrations on the resulting ceramic aerogel microstructures. The simulations were initialized with a fluid pool of height 400 $\mu m$, with bubbles of three distinct nuclei sizes (8 $\mu m$, 18 $\mu m$, and 25 $\mu m$) distributed within the simulation domain using a Poisson disk-sampling algorithm [35]. For all simulations, the fluid density was set to 1 $kg/{\mu m}^{3}$, ambient pressure to 0.33 $N/{\mu m}^{2}$, surface tension to $\sigma_{LB}=5\times 10^{-3}$ $N/\mu m$, and the advection-diffusion source term to $S=5\times 10^{-5}$ ${\mu m}^{2}/s$. The variation in the number of initial bubbles and their radii led to distinct classes of ceramic microstructures with different average pore size at the final simulation time. As demonstrated later, the use of the Poisson disk-sampling algorithm enables FSLBM to generate multiple ensemble realizations for each microstructure class with the same numbers and initial radius of the bubbles to account for stochasticity of the aerogel microstructure.
@@ 2.2 Generative surrogate model of aerogel microstructure
FSLBM simulations of aerogel microstructures are computationally intensive, making tasks such as uncertainty quantification, requiring numerous evaluations of $\mathcal{F}_{LB}(\boldsymbol{\xi};\boldsymbol{\theta}_{LB})$ prohibitively expensive. To enable computationally efficient synthesis-microstructure mapping, we propose a surrogate model for FSLBM simulations based on the Generative Adversarial Network (GAN) introduced by [36]. GANs employ two deep convolutional neural networks: the generator $\mathcal{G}_{GAN}(\cdot)$ and the discriminator $\mathcal{D}(\cdot)$. These networks are trained an adversarial manner in a two-player minimax game, where the generator, parameterized by $\boldsymbol{w}_{GAN}$, maps a random latent input vector to the data space $p_{G}$, approximating the real data distribution $p_{data}$. Meanwhile, the discriminator, with weight parameters $\boldsymbol{w}_{dis}$, acts as a classifier, outputting the probability of a sample being real or generated. In this unsupervised learning setup, the generator minimizes the divergence between $p_{G}$ and $p_{data}$, while the discriminator maximizes it, until both networks reach a Nash equilibrium. Once trained on microstructural imaging data, the generator can efficiently produce ensembles of microstructures that closely replicate the characteristics of the training data. In recent years, the application of GANs to imaging and digital material data has gained significant attention, demonstrating their potential in capturing realistic microstructural features, e.g., [37, 38, 39, 40, 41].
@@ 3 Physical and surrogate models of aerogel mechanical properties
This section introduces a physics-based model for analyzing stress and strain distribution within the two-phase aerogel material. Additionally, it describes the development of a microstructure-property surrogate model using a Bayesian convolutional neural network (BayesCNN). Coupled with the microstructural images generated by the WGAN-GP model outlined in Section 2.2, the linked-surrogates enables facilitates efficient prediction of the quantity of interest (QoI), specifically the strain energy.
@@ 3.1 Stochastic finite element model of aerogel elastic deformation
[…]
Figure 3: Finite element (FE) simulation results for three microstructure indicator functions, showing the magnitude of displacement and Cauchy stress fields. For the FE simulations, young’s modulus for solid $E_{s}=100\times 10^{6}$ Pa, poisson’s ratio $\nu$ = 0.3 was used. A traction force of $10^{6}$ N was applied in the negative y direction (compression).
The prediction QoI is the scalar strain energy, computed as:

### v3

@@ Abstract
This study presents an integrated computational framework that, given synthesis parameters, predicts the resulting microstructural morphology and mechanical response of ceramic aerogel porous materials by combining physics-based simulations with deep learning surrogate models. Lattice Boltzmann simulations are employed to model microstructure formation during material synthesis process, while a finite element model is used to compute the corresponding mechanical properties. To overcome the prohibitive computational demands of repeated physics-based simulations required for characterizing the impact of microstructure randomness on mechanical properties, surrogate models are developed using Convolutional Neural Networks (CNNs) for both microstructure generation and microstructure-property mapping. CNN training is formulated as a Bayesian inference problem to enable uncertainty quantification and provide confidence estimates in surrogate model predictions, under limited training data furnished by physics-based simulations. Numerical results demonstrate that the microstructure surrogate model effectively generates microstructural images consistent with the morphology of training data across larger domains. The Bayesian CNN surrogate accurately predicts strain energy for in-distribution microstructures and its generalization capability to interpolated morphologies are further investigated. Finally, the surrogate models are employed for efficient uncertainty propagation, quantifying the influence of microstructural variability on macroscopic mechanical property.
@@
Bayesian inference , convolutional neural networks , surrogate modeling , uncertainty propagation , ceramic aerogel
@@ 1 Introduction
Ceramic aerogels, composed of interconnected amorphous ceramic nanoparticles, are highly porous materials with exceptional thermal insulation properties, making them attractive for aerospace, microelectronics, fire-resistant wearables, and energy-efficient buildings [1, 2, 3, 4, 5]. However, their widespread adoption is limited by poor mechanical robustness. Physics-based synthesis-microstructure-property models offer a pathway to understanding and optimizing material performance, e.g., [6, 7, 8] to achieve mechanically robust aerogels without compromising insulation performance. Due to the inherent randomness (stochasticity) in microstructural features, such as pore size, shape, and spatial distribution, uncertainty propagation is essential for quantifying variability in mechanical properties. However, physics-based models are often computationally prohibitive for this task, as it demands large number of samples involving repeated model evaluations. To overcome this challenge, data-driven surrogate models [9, 10], trained on high-fidelity data from physics-based simulations, can efficiently approximate the underlying models, enabling uncertainty propagation at significantly reduced computational cost.
@@
Convolutional Neural Networks (CNNs) have become increasingly prominent in computational materials science [11, 12, 13, 14, 15, 16], particularly for generating realistic microstructural images and predicting macroscopic properties. However, deploying CNNs as surrogate models often requires training on sparse and limited datasets generated from computationally expensive physics-based simulations. While maximum likelihood estimation is the standard training approach, its performance can degrade in small-data or high-complexity regimes without appropriate safeguards. Overfitting and overconfident predictions may occur in such cases, but these risks are commonly mitigated through techniques such as regularization, cross-validation, transfer learning, and self-supervised learning [17, 18, 19, 20]. An alternative and complementary strategy is to formulate CNN training within a Bayesian inference framework [21, 22], where model parameters are treated as random variables with posterior distributions conditioned on the training data. In principle, Bayesian inference enables uncertainty quantification and provides confidence estimates in surrogate model predictions [23, 24]. However, the practical application of Bayesian CNNs in surrogate modeling remains limited due to the computational challenges posed by high-dimensional parameter spaces and the sensitivity to prior and likelihood specifications, both of which can significantly influence accuracy and reliability of model prediction. Zhu et al. [25] addressed the computational challenges by employing a variational gradient descent algorithm based on Stein’s method, extending Bayesian inference to convolutional encoder-decoder networks. Their approach captured uncertainty in flow through heterogeneous media, even in scenarios where the input (permeability) and output (flow and pressure) fields lacked shared underlying structures. Similarly, Shridhar et al. [26] applied variational inference to estimate posterior weight distributions in CNNs, effectively propagating uncertainty in classification tasks across datasets like MNIST and CIFAR-100.
@@
Figure 1: Flowchart of the integrated modeling framework developed in this work for propagating uncertainty through synthesis inputs $\boldsymbol{\xi}$, microstructure spatial field $\varphi(\mathbf{x})$, and mechanical property $Q$ of ceramic aerogels. The physics-based models, $\mathcal{F}_{LB}(\cdot)$ and $\mathcal{F}_{FE}(\cdot)$, are parameterized by $\boldsymbol{\theta}_{LB}$ and $\boldsymbol{\theta}_{FE}$, and utilize synthesis inputs and microstructural patterns to predict material properties. Computational efficiency is achieved through surrogate models $\mathcal{G}_{GAN}(\cdot)$ and $\mathcal{G}_{CNN}(\cdot)$, constructed using high-fidelity training data from physics-based simulations. The parameter of the surrogate models include the stochastic latent vector $\boldsymbol{z}$, deterministic weights of the generative network $\boldsymbol{w}_{GAN}$, and posterior probability distributions of the wights $\boldsymbol{w}_{CNN}$ from the Bayesian inference.
This work introduces an integrated modeling framework (Figure 1) for simulating the synthesis, microstructure, and properties of ceramic aerogels. Specifically, it focuses on a novel synthesis method utilizing in situ bubble-supported pore formation with foaming agents [27, 2], which addresses the limitations of conventional supercritical drying by preserving the porous structure without collapsing the solid matrix. This synthesis enables ambient-pressure, room-temperature drying, significantly reducing processing time while facilitating scalable and cost-effective production of ceramic aerogels. The synthesis-microstructure physics-based model in Figure 1, relies on Lattice Boltzmann simulations of the pre-aerogel foaming process, capturing stochastic bubble nucleation, bubble dynamics, and coalescence to predict microstructure formation. The microstructure-property model employs finite element simulations to analyze the elastic deformation of the solid phase of the aerogel. Using the microstructural images, this model computes the displacement and strain fields to estimate the mechanical properties of the aerogel, i.e., the strain energy as the quantity of interest (QoI) for the modeling framework. To enable efficient uncertainty propagation from microstructure to mechanical property, we develop surrogate models of the expensive physics-based simulations. The synthesis–microstructure surrogate employs a scalable generative adversarial network trained on Lattice Boltzmann-generated microstructures, allowing the generation of morphologically realistic microstructures across a range of domain sizes. The microstructure-property surrogate employs a CNN trained to map microstructural images to strain energy. A key innovation of this framework is the formulation of CNN training as Bayesian inference, enabling the estimation of the probability distribution of weight parameters using high-fidelity training data from finite element simulations. To address the computational challenges of Bayesian inference in high-dimensional parameter spaces, we extend the variational inference approach commonly used in Bayesian Neural Networks, e.g., [24] to accommodate the complexity of CNN architectures. The resulting Bayesian CNN (BayesCNN) enables not only prediction of the strain energy but also quantification of the associated uncertainty, providing a measure of confidence in the surrogate model’s predictions. These surrogates are then used to perform efficient uncertainty propagation analyses, that is, to quantify how microstructural randomness influences variability in mechanical properties. The generative model produces ensembles of statistically consistent microstructures, while the BayesCNN evaluates the corresponding strain energy to estimate the probability distribution and credible intervals of the QoI (strain energy) across microstructures with different average pore sizes.
@@
The remainder of this manuscript is structured as follows. Section 2 provides an overview of the ceramic aerogel synthesis process, including its modeling using Lattice Boltzmann simulations and the development of surrogate models to represent microstructural patterns. Section 3 describes the finite element model used to establish microstructure-property relationships and presents the comprehensive formulation and efficient solution algorithm of BayesCNN for surrogate modeling. Section 4 outlines the construction of individual surrogate models, detailing the creation of training datasets from high-fidelity physics-based simulations and the validation of these models. Additionally, this section illustrates the use of the surrogate models to efficiently propagate uncertainty from stochastic microstructure to the predicted QoI. Finally, Section 5 presents a discussion of the findings and the concluding remarks.
@@ 2.1 Lattice Boltzmann simulation of microstructure formation
where $\Delta t$ is the time step, $\tau$ is a relaxation time, and $f_{i}^{eq}$ is the distribution function at equilibrium [37]. The solution of Eq. (1) involves two iterative steps: streaming, where the distribution functions $f_{i}$ move to neighboring lattice sites, and collision, where $f_{i}$ is locally updated at each site. Macroscopic fluid properties, such as density and velocity, are then derived from the moments of $f_{i}$. Building on this foundation, FSLBM simulations utilize a fast mass-tracking algorithm [31] to monitor the interfaces between gas and liquid phases throughout the computational domain. The liquid-gas interface is modeled as a free surface, with the Volume-of-Fluid (VOF) method employed to track the interface using a level function, which represents the liquid fraction within each computational cell. Surface tension $\sigma_{LB}$ are incorporated by calculating the interface curvature from the level function and applying the resulting pressure jump across the interface. In this work, to capture the highly viscous and shear-thinning behavior of the aerogel precursor, governed by the concentration of viscosifier $C_{vis}$, the Carreau-Yasuda model [38] is adopted to describe the fluid’s rheological properties. The model provides an accurate representation of the non-Newtonian characteristics essential for simulating the foaming process. The influence of TEOS concentration on bubble growth is incorporated by modeling gas diffusion into the bubbles and within the liquid phase using the advection-diffusion equation for gas concentration $C$,
@@
Figure 2 presents snapshots of pre-aerogel foaming simulations conducted using FSLBM. The simulations were initialized with a fluid pool of height 400 $\mu m$, with bubbles of three distinct nuclei sizes (8 $\mu m$, 18 $\mu m$, and 25 $\mu m$) spatially distributed within the simulation domain using a Poisson disk-sampling algorithm [39]. The initial number and radii of bubbles serve as proxies for varying concentrations of the foaming agent, one of the key synthesis parameters influencing the resulting ceramic aerogel microstructure. For all simulations, the fluid density was set to 1 $kg/{\mu m}^{3}$, ambient pressure to 0.33 $N/{\mu m}^{2}$, surface tension to $\sigma_{LB}=5\times 10^{-3}$ $N/\mu m$, and the advection-diffusion source term to $S=5\times 10^{-5}$ ${\mu m}^{2}/s$. We note that, the FSLBM framework accounts for bubble dissolution by coupling the Lattice Boltzmann flow solver with an advection-diffusion equation subject to Henry’s law boundary conditions at bubble interfaces. Bubble rupture is modeled using a disjoining-pressure formulation, where coalescence or collapse occurs if the stabilizing forces between interfaces are insufficient. However, in the ceramic aerogel synthesis process, the combination of CTAB micelles and TEOS-driven silica condensation rapidly forms rigid shells around nascent bubbles, while urea decomposition introduces gas into a gelling silica matrix. This coordinated mechanism effectively stabilizes bubbles against rupture and dissolution, an effect captured by the parameters used in our FSLBM simulations. The variation in the number of initial bubbles and their radii led to distinct classes of ceramic microstructures with different average pore size at the final simulation time. As demonstrated later, the use of the Poisson disk-sampling algorithm enables FSLBM to generate multiple ensemble realizations for each microstructure class with the same numbers and initial radius of the bubbles to account for stochasticity of the aerogel microstructure.
@@ 2.2 Generative surrogate model of aerogel microstructure
FSLBM simulations of aerogel microstructures are computationally expensive, particularly for tasks involving uncertainty propagation through the microstructure–property relationship (see Figure 1), which require repeated evaluations of $\mathcal{F}_{LB}(\boldsymbol{\xi};\boldsymbol{\theta}_{LB})$ to capture the inherent stochasticity of the microstructure. To enable a computationally efficient synthesis-to-microstructure mapping, we introduce a surrogate model for the FSLBM simulations based on the Generative Adversarial Network (GAN) framework originally proposed by [40]. GANs employ two deep convolutional neural networks: the generator $\mathcal{G}_{GAN}(\cdot)$ and the discriminator $\mathcal{D}(\cdot)$. These networks are trained an adversarial manner in a two-player minimax game, where the generator, parameterized by $\boldsymbol{w}_{GAN}$, maps a random latent input vector to the data space $p_{G}$, approximating the real data distribution $p_{data}$. Meanwhile, the discriminator, with weight parameters $\boldsymbol{w}_{dis}$, acts as a classifier, outputting the probability of a sample being real or generated. In this unsupervised learning setup, the generator minimizes the divergence between $p_{G}$ and $p_{data}$, while the discriminator maximizes it, until both networks reach a Nash equilibrium. Once trained on microstructural imaging data, the generator can efficiently produce ensembles of microstructures that closely replicate the characteristics of the training data. In recent years, the application of GANs to imaging and digital material data has gained significant attention, demonstrating their potential in capturing realistic microstructural features, e.g., [41, 42, 43, 44, 45].
@@ 3 Physical and surrogate models of aerogel mechanical properties
This section introduces a physics-based model for analyzing stress and strain distribution within the two-phase aerogel material. Additionally, it describes the development of a microstructure-property surrogate model using a Bayesian convolutional neural network (BayesCNN). Coupled with the microstructural images generated by the WGAN-GP model outlined in Section 2.2, the surrogate models facilitate efficient prediction of the probability distribution of the quantity of interest (QoI), specifically the strain energy.
@@ 3.1 Stochastic finite element model of aerogel elastic deformation
where, $\lambda_{s}$ and $\mu_{s}$ are the Lamé constants of the solid aerogel phase, equivalent to the material’s Young’s modulus $E_{s}$ and Poisson’s ratio $\nu_{s}$, $\boldsymbol{E}(\boldsymbol{u})=\frac{1}{2}\left(\nabla\boldsymbol{u}+(\nabla\boldsymbol{u})^{T}\right)$ is the strain tensor, and $\varphi(\mathbf{x})$ is the stochastic microstructure indicator function generated by the WGAN-GP surrogate model described in the previous section. Figure 3 illustrates the displacement and stress fields computed for a single realization of aerogel microstructure. These results are obtained under uniaxial compression conditions, where a traction force of $t_{y}=10^{5}$ N is applied in the negative $y$-direction at the top surface. The bottom boundary is fixed $\boldsymbol{u}^{*}=0$, while roller boundary conditions $u^{*}_{x}=0$ are imposed at the left and right boundaries. The governing equation in (3.1) is solved using a uniformly fine mesh to resolve the intricate microstructural patterns represented by $\varphi(\mathbf{x})$. Convergence analysis with respect to mesh resolution indicates that a finite element mesh with 358$\times$358 elements provides sufficient accuracy and numerical stability for this class of microstructures (see Figure 3(b)). All finite element simulations are carried out using the DOLFINx library [48] from the FEniCS Project in Python. The prediction QoI is the scalar strain energy, computed as:
@@
Figure 3: Finite element (FE) simulation results for for a representative microstructure indicator function: (a and b) Microstructure image and mesh convergence study comparing two mesh resolutions, 720 $\times$ 720 and 358 $\times$ 358 elements, yielding strain energies of $1382\times 10^{-7}$ mJ and $1378\times 10^{-7}$ mJ, respectively. The relative error of 0.3% confirms that the 358 $\times$ 358 mesh provides sufficient resolution for accurate results and is used for all subsequent simulations. (c and d) Magnitude of the computed displacement (in mm) and Cauchy stress (in mPa=$10^{-3}N/m^{2}$) fields under uniaxial compression. For all FE simulations, the solid phase is assigned a Young’s modulus $E_{s}=100\times 10^{6}$ Pa and Poisson’s ratio $\nu_{s}$ = 0.3. A traction force of $10^{5}$ N was applied in the negative y-direction.
We note that in this study, a linear elastic constitutive model is adopted for the solid phase of the ceramic aerogel, to demonstrate the proposed surrogate modeling framework. However, this assumption may not capture nonlinear deformation mechanisms or damage evolution that may arise under loading and ultimately lead to mechanical failure. Ongoing mechanical testing of ceramic aerogels fabricated via in situ bubble-supported synthesis would better inform the development of more representative constitutive models that incorporate these effects (e.g., [27, 49, 50]).